This frequency distribution shows the number of pounds of each snack food eaten during the Super Bowl. Construct a pie graph for the data (20 points). Snacks Potato chips Tortilla chips Pretzels Pop Corn Snack nuts Pounds (Frequency) 11 million 8 million 4 million 3 million 2 million a) What is mode b) What is the sample size c) From the table above, construct Categorical Frequency distribution showing the relative frequencies (percent) column. d) Construct a Pareto Chart for this data ensuring that all the labels and title are clearly stated.
Answer #1

Sacks unds 3mackP 36o 28 Potato chips | il millon|--x360 Pketzels Snack nuts | 2 rìnrot 읊 x 360 、41-42~141 2. 8 Totat 2 8million 360 เค. Palal, cups 々60 t. Snack nuts 103 37 Chips Si opcon

a) Since the highest frequency corresponds to Potato Chips, that is 11 million the mode of the given data is Potato Chips.

b) Total of the frequency=(11+8+4+3+2)million=28 million=sample size

c)

wnds Frqueney Potato chips l millon 23 Psetzels 4 milluow PopCom mill on rack nut2 mill oir 2.3 4 X 100 = 14-285 28 26 xIoO 10-1142 28 28 2 28

d)Pareto Chart: A chart which is represented by bars of frequency in descending order and line graph of cumulative frequencies